# SAGMAN: Stability Analysis of Graph Neural Networks on the Manifolds

### v3

@@ Appendix C SAGMAN without GDR
Appendix D Fast Effective-Resistance Estimation for LRD Graph Decomposition
@@
Appendix E Additional Results for GNN Stability Evaluation and Statistics of Datasets
We present the additional results in Figure 8, Figure 9, Figure 10, and Table 7. Table 6 summarizes the datasets utilized.
@@ Appendix E Additional Results for GNN Stability Evaluation and Statistics of Datasets
Figure 8: Figures represent cosine similarities and KL divergence. ”Random EdgesPT” corresponds to the DICE edge evasion attack. ”Random FeaturePT” refers to Gaussian noise evasion perturbation $X+\xi\eta$, where $X$ is feature matrix, $\eta$ is Gaussian noise, $\xi$ is noise level controls
@@
Appendix F Why Generalized Eigenpairs Associate with DMD
Cheng et al. propose a method to estimate the maximum distance mapping distortion (DMD), denoted as $\delta^{M}_{max}$, by solving the following combinatorial optimization problem:
@@ Appendix F Why Generalized Eigenpairs Associate with DMD
In this study, we restrict our calculations of stability scores to the two largest generalized eigenpairs. We also show results of the third and fourth largest generalized eigenpairs, as well as the fifth and sixth largest generalized eigenpairs in Table 8.
Table 8: Cosine Similarity and KL Divergence for DICE Edges Perturbation and Random Feature Perturbation. We bold the better results.
Largest Generalized Eigenpairs: 1st and 2nd/3rd and 4th/5th and 6th
Method PT Level Cosine KL Divergence
Stable Unstable Stable Unstable
EdgesPT
1
0.99/0.98/0.99 0.97/0.96/0.91 0.04/0.25/0.12 0.46/0.73/1.83
EdgesPT
2
0.96/0.87/0.96 0.92/0.89/0.86 0.82/1.17/1.11 2.43/2.01/3.47
EdgesPT
3
0.88/0.81/0.90 0.92/0.87/0.84 2.18/2.29/2.04 2.71/3.79/5.34
EdgesPT
4
0.83/0.80/0.85 0.88/0.85/0.85 4.57/3.29/3.83 4.16/4.23/5.25
EdgesPT
5
0.81/0.78/0.82 0.88/0.83/0.76 3.88/4.56/5.28 4.92/4.07/6.02
FeaturePT
0.4
1.00/0.84/0.81 1.00/0.75/0.68 0.00/2.16/5.70 0.00/4.57/8.24
FeaturePT
0.8
1.00/0.79/0.75 1.00/0.73/0.71 0.00/4.00/6.49 0.01/4.87/8.29
FeaturePT
1.2
0.98/0.78/0.76 0.98/0.70/0.69 0.02/6.11/6.65 0.08/5.47/9.33
FeaturePT
1.6
0.97/0.69/0.71 0.93/0.70/0.64 0.01/8.94/8.85 1.09/7.07/8.41
FeaturePT
2.0
0.80/0.67/0.70 0.82/0.72/0.66 1.72/8.57/9.53 4.02/5.45/9.40
Appendix G Various Sampling Schemes for SAGMAN-guided Perturbations
Previous works (Cheng et al., 2021; Hua et al., 2021; Chang et al., 2017) highlighted that only part of the dataset plays a crucial role in model stability, so we want to focus on the difference between the most ”stable” and ”unstable” parts. However, it is certainly feasible to evaluate the entire graph. Table 9 shows the result regarding the Pubmed dataset in GPRGNN under Gaussian noise perturbation. Samples were segmented based on SAGMAN ranking, with the bottom 20% being the most ”stable”, the middle 60% as intermediate, and the top 20% representing the most ”unstable”. As anticipated, the ”stable” category (representing the bottom 20%) should exhibit the lowest average KL divergences. This is followed by the intermediate category (covering the mid 60%), and finally, the ”unstable” category (comprising the top 20%) should display the highest divergences.

### v4

@@ A.8 GNN Stability Analysis without GDR
A.9 Fast Effective-Resistance Estimation for LRD-based Graph Decomposition
@@
A.10 Additional Results for GNN Stability Evaluation and Statistics of Datasets
We present the additional results in Figure 8, Figure 9, Figure 10, Figure 7, and Table 9. Table 8 summarizes the datasets utilized.
Figure 7: SAGMAN-selected stable/unstable samples for the "ogbn-arxiv" dataset. We report the GCN prediction accuracies under different levels of Gaussian Noise Perturbation $FM+X\eta$, where $FM$ denotes the feature matrix, $\eta$ represents Gaussian noise, and $X$ is the noise perturbation level.
@@ A.10 Additional Results for GNN Stability Evaluation and Statistics of Datasets
Figure 8: Figures represent cosine similarities and KL divergence. "Random EdgesPT" corresponds to the DICE edge evasion attack. "Random FeaturePT" refers to Gaussian noise evasion perturbation $X+\xi\eta$, where $X$ is feature matrix, $\eta$ is Gaussian noise, $\xi$ is noise level controls
@@
A.11 Why Generalized Eigenpairs Associate with DMD
(Cheng et al., 2021) propose a method to estimate the maximum distance mapping distortion (DMD), denoted as $\delta^{M}_{max}$, by solving the following combinatorial optimization problem:
@@ A.11 Why Generalized Eigenpairs Associate with DMD
A.12 Various Sampling Schemes for SAGMAN-guided Perturbations
Previous works (Cheng et al., 2021; Hua et al., 2021; Chang et al., 2017) highlighted that only part of the dataset plays a crucial role in model stability, so we want to focus on the difference between the most "stable" and "unstable" parts. However, it is certainly feasible to evaluate the entire graph. Table 10 shows the result regarding the Pubmed dataset in GPRGNN under Gaussian noise perturbation. Samples were segmented based on SAGMAN ranking, with the bottom 20% being the most "stable", the middle 60% as intermediate, and the top 20% representing the most "unstable". As anticipated, the "stable" category (representing the bottom 20%) should exhibit the lowest average KL divergences. This is followed by the intermediate category (covering the mid 60%), and finally, the "unstable" category (comprising the top 20%) should display the highest divergences.
@@ A.12 Various Sampling Schemes for SAGMAN-guided Perturbations
A.13 SAGMAN’s Runtime Across Datasets
Figure 11 demonstrates the high efficiency of SAGMAN in processing large graphs, attributing its performance to the near-linear time complexity of its components.
Figure 11: SAGMAN’s runtime across datasets